10 partners agree not to have the assets revalued upon Bensons retirement nterest the eputy. Pre10 of 12 Next> O Type here to search
《 Prey 10:of12 Type here to search
on Rhode Invests (a) $116,667: (b) $ 《 Prey 0 of 12 Type here to search
Answer #1

1.

Feb. 1 Cash $116,667
Rhodes' Capital $116,667

$33,000+$139,000+$178,000= $350,000

($350,000+$116,667)*25% = $116,667

No bonus paid

2.

Feb. 1 Cash $85,167
Meir capital ($23,625*1/10) $2,362
Benson capital($23,625*4/10) $9,450
Lau capital($23,625*5/10) $11,813
Rhode capital $108,792

($350,000+$85,167)*25% = $108,792

($108,792 - $85,167) = $23,625

$23,625 bonus paid to new partner.

3.

Feb. 1 Cash $152,834
Meir capital($27,126*1/10) $2,713
Benson capital ($27,126*4/10) $10,850
Lau capital ($27,126*5/10) 13,563
Rhode capital $125,708

($350,000+$152,834)*25% = $125,708

($152,834 - $125,708) =$27,126

$27,126 bonus to old partners in ratio 1:4:5.